Given the demand function Qx = 8000-1000Px, determine the elasticity of demand at a single point where price of Rs. 6 and the corresponding quantity of 2000 units


1
Expert's answer
2022-03-24T16:12:38-0400


Qx = 8000-1000Px


Ped="\frac {percentage change in Q}{percentage change in P}"


Ped="\frac{dq}{dp}\times\frac{p}{q}" ,where p=6,q=2000


"\frac{dq}{dp}" 8000-1000Px=-1000 therefore


=-1000"\times" "\frac{6}{2000}"


="\frac{-6000}{2000}" =-3


elasticity of demand is "\mid"-3"\mid" ,hence

PEd=3